# Exploring Quantization for Efficient Pre-Training of Transformer Language Models

### v1

@@ 2 Related Work
In contrast, Fully Quantized Training (FQT) harnesses the accelerated gains derived from higher throughput in INT8 or INT4 operations supported by modern GPUs during training. Additionally, FQT capitalizes on memory savings by storing parameters in lower precisions. As exemplified by Li et al. (2024) and Dettmers et al. (2021), states of the Adam optimizer are stored in 4 and 8 bits, respectively, to minimize memory footprint. Another strategy introduced by Markov et al. (2023) involves quantizing both weights and gradients to reduce bandwidth usage in distributed training. Furthermore, Wortsman et al. (2024) and Kim et al. (2021) advocate the replacement of linear operations with INT8 matrix multiplications to achieve substantial speedups.
@@
Wortsman et al. (2024) employ 8-bit quantization for linear operations in both forward and backward passes, achieving remarkable results in pre-training large-scale vision language models. Their approach incorporates row-wise quantization for activations and gradients, mitigating the impact of quantization errors on other parameters. However, vision language models significantly differ from large textual language models. Xi et al. (2024b) proposed 4-bit quantization in FQT with Hadamard transformation to address outliers in activations. Acknowledging the sparsity of gradients during training, they propose bit splitting for quantizing gradients in 4-bit precision. However, they predominantly focus on fine-tuning, and their experiments on pre-training only consider small models (60M) on a limited dataset (WMT 14 En-De).
@@
In contrast to the aforementioned complex quantization methods, such as quantile quantization and learnable quantization parameters, this study focuses on a more straightforward implementation. The objective is to investigate the effects of quantization and explore the feasibility of training LLMs with full integer operations. The primary goal of this paper is to offer detailed insights into quantizing different components of the model. While more sophisticated quantization methods have the potential to enhance the final performance, our work serves as a foundational investigation, providing valuable insights and paving the way for further exploration in the realm of quantization for LLMs.
@@ 3.1 Quantization Scheme
Figure 1 illustrates how quantization error is introduced during both the forward and backward passes. Specifically, the quantization error is injected into either the weights or input activations during the forward pass. In the backward pass, quantization is applied to the output gradient for computing weight updates due to only the gradient of the weight is stored. Nonetheless, the real-valued output gradient is used to compute the quantized input gradient since we observed an increase in training instability when propagating the quantization error through the entire backward path. Additional details about gradient quantization are presented in Section 4.3.
@@ 4 Experimental Results
We used GPT-2 small (Radford et al., 2019) (124M) with FlashAttention2 (Dao et al., 2022) for our experiments due to its popularity and remarkable performance-to-compute ratio. For the following experiments, we pre-trained 30 models from scratch on OpenWebText (Gokaslan and Cohen, 2019) for 300k gradient steps with a global batch size of 512 samples and a context length of 1024 tokens. For our evaluation setup, we evaluate the performance of the models on a range of language tasks, including ARC-Easy (Yadav et al., 2019), ARC-Challenge (Yadav et al., 2019), Hellaswag (Zellers et al., 2019), LAMBADA (Paperno et al., 2016), and GLUE score (Wang et al., 2018). Additional details about our training and evaluation setups are provided in Appendix A.
@@ Appendix B Memory Analysis
Figure 15: Peak memory usage profile for different sequence lengths while maintaining a constant batch size of 4, across different model sizes (GPT-2 Small, Medium, and Large). This figure illustrates the shift in peak memory usage from gradients to activations as sequence length increases, and the significant impact of sequence length on memory dynamics during training.
In this section, we explore the memory consumption patterns of various components within GPT-2 models during training, using the PyTorch Memory Profiler. This profiling tool allows for precise monitoring of memory usage throughout the lifecycle of a model’s operation, particularly during the forward and backward passes, and during optimization steps. "Peak memory" in this context refers to the maximum memory usage observed at any point in these stages, providing insights into how different model configurations impact overall memory requirements. We profiled all attention blocks as well as lm head of the Transformer model.
Figure 14 shows how memory usage changes with different batch sizes for a fixed context length of 1024, across various model sizes. Notably, as the batch size increases, the memory allocated for activations becomes more dominant, especially in larger models. This trend is primarily due to the need to store activations for the computation of gradients during the backward pass, which increases with larger batch sizes.
Figure 15 examines the memory usage across various sequence lengths while keeping the batch size constant at 4. When both batch size and sequence length are small, peak memory typically occurs towards the end of the backward propagation phase. At this stage, memory includes the parameters, optimizer states, gradients from all layers, and activations from the initial layers. However, as the sequence length and batch size increase, peak memory usage shifts to the beginning of the backward propagation. At this point, the memory comprises parameters, optimizer states, all activations, and notably, the output gradient of the final layers, which matches the size of the logits (proportional to batch size * sequence length * vocabulary size).
@@ Appendix C Quantization Results
This appendix presents the granular results from our experiments on weight, activation, gradient, and optimizer states quantization. The tables detail the performance metrics, like perplexity and accuracy, under various quantization settings. Finally, we present post-training quantization results of our baseline model in Tables 10,11.
Table 10: Post-training weight quantization results
@@
Table 11: Post-training activation quantization results

### v2

@@ 2 Related Work
In contrast, Fully Quantized Training (FQT) harnesses the accelerated gains derived from higher throughput in INT8 or INT4 operations supported by modern GPUs during training. Additionally, FQT capitalizes on memory savings by storing parameters in lower precisions. As exemplified by Li et al. (2024) and Dettmers et al. (2021), states of the Adam optimizer are stored in 4 and 8 bits, respectively, to minimize memory footprint. Another strategy introduced by Markov et al. (2023) involves quantizing both weights and gradients to reduce bandwidth usage in distributed training. Furthermore, Wortsman et al. (2024) and Kim et al. (2021) advocate the replacement of linear operations with INT8 matrix multiplications to achieve substantial speedups. Our work extends quantization to multiple components, providing a more comprehensive exploration of the challenges and benefits of quantization during pre-training.
@@
Wortsman et al. (2024) employ 8-bit quantization for linear operations in both forward and backward passes, achieving remarkable results in pre-training large-scale vision language models. Their approach incorporates row-wise quantization for activations and gradients, mitigating the impact of quantization errors on other parameters. However, vision language models significantly differ from large textual language models. Xi et al. (2024b) explored 4-bit quantization in Fully Quantized Training (FQT) using Hadamard transformations to handle activation outliers and proposed bit splitting to quantize gradients in 4-bit precision. However, their work primarily focuses on fine-tuning, and only relatively small models (60M parameters) are pre-trained on a small dataset (WMT 14 En-De), which hinders the generalizability of their findings to larger language models. Our study explores pre-training at a larger scale, utilizing a bigger model and dataset to thoroughly examine the effects of quantization on various model components.
@@
In contrast to the aforementioned complex quantization methods, such as quantile quantization and learnable quantization parameters, this study focuses on a more straightforward implementation. The objective is to investigate the effects of quantization and explore the feasibility of training LLMs with full integer operations. The primary goal is to offer detailed insights into quantizing different components of the model. While more sophisticated quantization methods may enhance performance, our work serves as a foundational investigation, providing insights and paving the way for further exploration in quantization for LLMs.
@@ 3.1 Quantization Scheme
Figure 1 illustrates how quantization error is introduced during both the forward and backward passes. Specifically, the quantization error is injected into either the weights or input activations during the forward pass. Since we only store the gradients related to the weights in memory, the quantization in the backward pass is only applied to the output gradients (“grad out” in Figure 1) for weight updates. Nonetheless, the real-valued output gradient is used to compute the quantized input gradient since we observed an increase in training instability when propagating the quantization error through the entire backward path. Additional details about gradient quantization are presented in Section 4.3.
@@ 4 Experimental Results
We used GPT-2 small (124M) (Radford et al., 2019) with FlashAttention-2 (Dao et al., 2022) for our experiments due to its popularity as a baseline architecture for contemporary studies of LLMs. While larger models are often used, GPT-2 small provides a manageable framework for in-depth experimentation without requiring the immense computational resources of larger models. For our experiments, we pre-trained 30 models from scratch on OpenWebText (Gokaslan and Cohen, 2019) for 300k gradient steps with a global batch size of 512 samples and a context length of 1024 tokens, processing approximately 157 billion tokens, which is consistent with similar works (Liu et al., 2023; Dao et al., 2022). For our evaluation setup, we evaluate the performance of the models on a range of language tasks, including ARC-Easy (Yadav et al., 2019), ARC-Challenge (Yadav et al., 2019), Hellaswag (Zellers et al., 2019), LAMBADA (Paperno et al., 2016), and GLUE (Wang et al., 2018). Additional details about our training and evaluation setups are provided in Appendix A.
@@ Appendix B Memory Analysis
Figure 15: Peak memory usage profile for different sequence lengths while maintaining a constant batch size of 4, across different model sizes (GPT-2 Small, Medium, and Large). This figure illustrates the shift in peak memory usage from gradients to activations as sequence length increases and the significant impact of sequence length on memory dynamics during training.
In this section, we explore the memory consumption patterns of various components within GPT-2 models during training using the PyTorch Memory Profiler. This profiling tool allows for precise monitoring of memory usage throughout the lifecycle of a model’s operation, particularly during the forward and backward passes and during optimization steps. "Peak memory" in this context refers to the maximum memory usage observed at any point in these stages, providing insights into how different model configurations impact overall memory requirements. We profiled all attention blocks as well as the LM head of the Transformer model.
Figure 14 shows how memory usage changes with different batch sizes for a fixed context length of 1024 across various model sizes. Notably, as the batch size increases, the memory allocated for activations becomes more dominant, especially in larger models. This trend is primarily due to the need to store activations for the computation of gradients during the backward pass, which increases with larger batch sizes.
Figure 15 examines the memory usage across various sequence lengths while keeping the batch size constant at 4. When both batch size and sequence length are small, peak memory typically occurs towards the end of the backward propagation phase. At this stage, memory includes the parameters, optimizer states, gradients from all layers, and activations from the initial layers. However, as the sequence length and batch size increase, peak memory usage shifts to the beginning of the backward propagation. At this point, the memory comprises parameters, optimizer states, all activations, and notably the output gradient of the final layers, which matches the size of the logits (proportional to batch size * sequence length * vocabulary size).
@@ Appendix C Quantization Results
This appendix presents the granular results from our experiments on weight, activation, gradient, and optimizer states quantization. The tables detail the performance metrics, like perplexity and accuracy, under various quantization settings. Finally, we present post-training quantization results of our baseline model in Tables 10 and 11.
Table 10: Post-training weight quantization results.
@@
Table 11: Post-training activation quantization results.